OpenStudy (anonymous):

Please help! Use the formulas for functions of the difference and sum of two angles to find tan15degrees (using 45degrees and 30degrees) a. (sqrt3-3)/(sqrt3+3) b. (3-sqrt3)/(3+sqrt3) c. (3+sqrt3)/(3-sqrt3) d. (sqrt3+3)/(sqrt3-3)

OpenStudy (anonymous):

tan45-tan15

OpenStudy (anonymous):

tan45-tan30

OpenStudy (dumbcow):

\[\tan(15) = \tan(45-30) = \frac{\tan(45) -\tan(30)}{1+\tan(45)\tan(30)}\]

OpenStudy (dumbcow):

tan(45) = 1 tan(30) = sin(30)/cos(30) = (1/2) / sqrt3/2 = 1/sqrt3

OpenStudy (dumbcow):

did that help...do you know the answer now

OpenStudy (anonymous):

I got a. Can you please clarify? @dumbcow

OpenStudy (dumbcow):

it can't be A because tan(15) is positive both options A and D give negative answers i used the difference formula for tangent function i obtained values for tan(45) and tan(30) from unit circle then just plug them into formula and simplify it into single fraction Note: to get the answer they want....rationalize the denominator \[\frac{1}{\sqrt{3}} = \frac{\sqrt{3}}{3}\] \[\rightarrow \frac{1-\frac{\sqrt{3}}{3}}{1+\frac{\sqrt{3}}{3}}\]

OpenStudy (anonymous):

b.! THANKS SO MUCH